OpenStudy (anonymous):

Solve for y in the equation Y/5 =-15 Is y=-75 right? just checking as i go if thats ok

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thank you

OpenStudy (angela210793):

multiply with a 5 both sides y=-15*5=-75

OpenStudy (anonymous):

Solve for y in the equation y/2 =13 is y=26 right??

OpenStudy (angela210793):

yea

OpenStudy (anonymous):

yup :)

OpenStudy (anonymous):

Yes :P

OpenStudy (anonymous):

O crap i got something hard to do after this one brb -_-